OpenStudy (anonymous):

Find the inverse of y=6x-2

OpenStudy (anonymous):

*inverse equation

OpenStudy (goldphenoix):

\[\large \large y=6x-2\] So to find the inverse of it, you want x to be the only number in one of the side. To do this, you want to do opposite PEMDAS. So let's get rid of -2! To get rid of -2, you want to add both side by 2. -2 + 2 would give you 0. (Cancel out) y + 2 is equal to y + 2. (Since we don't know the variable) Your equation should look like: \[\large \large y + 2 = 6x\] The final step is to get rid of 6. The only way to get rid of a coefficent (a number times a variable) is to divide by the number. So 6x divided by 6 is x. y + 2 divided by 6 is y+2 divided by 6. So your equation should finaly be: \[\large \large\frac{ y +2 }{ 6 } = x\] I hope this help!

OpenStudy (goldphenoix):

Let me check my work.

OpenStudy (goldphenoix):

Not sure if this is inverse equation.